OpenStudy (anonymous):

Rewrite the equation in Standard Form. y = 2x + 5

jhonyy9 (jhonyy9):

this is the standard form of

OpenStudy (anonymous):

That's what it asked me to do?

jhonyy9 (jhonyy9):

so i think you havent what to do for

jhonyy9 (jhonyy9):

nothing

jhonyy9 (jhonyy9):

in this case

jhonyy9 (jhonyy9):

because the standard form is y=ax+b

OpenStudy (anonymous):

It gave me these options : A. -2x + y = 5 B. 2x - y = 5 C. 2x - y = -5 D. x - 1/2y = -5/2

jhonyy9 (jhonyy9):

the standard form is how have wrote there